Planning week 19-26 June 2010

The week began with AR 11082 being the only active region on the Sun, close to the central meridian. A couple of old, decaying active regions started coming around the east limb on 21 June. A nice equatorial coronal hole passed by the central meridian on 23-24 June.

The week was divided into four OP periods: OP1 was 2 days long, OP2 and OP3 were 2 days long, and OP4 was 1 day long. (These non-standard OP periods are due to a clash with the Akari satellite's telemetry plans.)

Note that this planning week took place during the Hinode eclipse season. Where possible studies were only scheduled for the daytime periods of the orbits.

Key observations

  1. HOP 150 (erupting filament): 19 Jun 11 UT - 20 Jun 10 UT.
  2. Bright point trigger test: 20 Jun and 24 Jun.
  3. Klimchuk nanoflare study: 21 Jun, 11 UT.
  4. AR spectral atlas: 21 Jun, 24 Jun.
  5. HOP 162 on active region: 22 Jun.
  6. HOP 73 (quiescent prominence): 22 Jun.
  7. HOP 161 (AR loop oscilations): 22-23 Jun, east limb.

Saturday 19 June (11-24 UT)

AIA 193 image at 22:09 UT
GOES plot (19-21 Jan)

A nice filament was present close to AR 11082 on the west side of the AR so the ToO HOP 150 (erupting filaments) was called. EIS ran the study AR_filament from around 11 UT to around 10 UT the next day.

Sunday 20 June (00-24 UT) 

AIA 193 image at 23:24 UT

After completing support of HOP 150, EIS performed the first ever test of the bright point trigger at 10:40 UT at disk center. No more studies were performed this day. UPDATE: this test was successful.

Inspection of AIA images suggested the morphology of AR 11082 changed significantly during 20 June due to the emergence of a nearby bright point. The HOP 150 data may thus show something interesting.

Monday 21 June (00-11 UT)

AIA 193 image at 23:28 UT

A test of the new study HPW021 was inserted at 01:15 and pointed at the east footpoints of the active region. No other science studies were run in this OP.

Monday 21 June (11-24 UT)

GOES plot, 21-23 Jun

I ran the nanoflare study at 11:06 and 12:45 on the east and west footpoints of AR 11082, respectively. The latter was repeated twice.

The spectral atlas study, Atlas60, was run on AR 11082 at 14:23. No more science studies were run this day.

Tuesday 22 June (00-24 UT)

AIA 193 image at 20:39 UT

As part of HOP 162 (solar atmosphere from c/s to corona), a SUMER HOP coordinated by Harry, I ran HPW021 at 00:14 and 01:53. The pointings are side-by-side and cover most of AR 11082.

HOP 73 (quiescent prominence) was supported from 13:19 to 17:13 with runs of prom_rast_v1, BL_SUMER_EIS_p2 and PRY_slot_context_v3. The target was a prominence at the NW limb, and AIA images show that the prominence was definitely there at the time of observation. UPDATE: the EIS raster was positioned too far to the north so the prominence appears at the bottom of the 195 image and is barely seen in 256.

HOP 161 (AR oscillations) was supported from 23:13 to 05:05 UT the next day via sit-and-stare observations with DEI_AR_LOOP_V2 over 4 orbits. The pointing was the old active region appearing at the east limb.

Wednesday 23 June

AIA 193 image
GOES plot, 23-25 Jun

HOP 161 was supported until 05:05 UT - see above.

There's a nice equatorial coronal hole and so HOP 163 (Louise's coronal hole study) was run from 12:20 to 16:00. I took the opportunity to run a couple of extra studies: SK_DEEP_10x512_SLIT2 was run using ASRC to make a 100x512 raster (7 spaced pointings) which might be useful for velocity studies. I also slipped in a Atlas_60s partly in the coronal hole and partly in the quiet Sun.

Thursday 24 June

AIA 193 image

The bright point trigger test is performed again from 03:15. The pointing is just north of disk center to avoid the equatorial coronal hole. UPDATE: this test was successful.

HOP 162 is supported on AR 11082 from 09:40 to 15:45. Three runs of HPW021 are placed adjacently across the region, with Atlas_60 run afterwards to obtain a spectral atlas.

Friday 25 June

AIA 193 image
GOES plot, 25-26 Jun

HOP 146 (Kuen's coronal hole monitoring) is supported from 02:00 to 07:00 at the south pole. The coronal hole is quite small.

HOP 73 (quiescent prominence) is supported again from 13:30 to 16:30 (two orbits) at the NW limb, i.e., same prominence as last time.

HOP 162 is supported at disk center from 18:30 to 21:30 (two orbits).

The bright point trigger test is performed again at disk center from 23:30.

Saturday 26 June (00-11 UT)

AIA 193 image

The only study in this period is the bright point trigger test from 23:30 to 01:00.